OpenStudy (anonymous):

HEEEELP :c What is the area of a sector with measure of arc equal to 90° and radius equal to 1 foot? answer should be Area= _________pi sq. ft. :(

OpenStudy (anonymous):

area of the circle: A=pi*R^2 sector 90 degree = 1/4 of the circle (circle = 360), right? then you area/ area of the sector: A=1/4 pi*1=1/4 pi

OpenStudy (anonymous):

pi/4

OpenStudy (anonymous):

OMG:) its right!!! thats the first time this openstudy thing has worked for me!! thank you soooo much!!! im doing this stuff online and one of the many questions that doesnt let me pass it is that one!! thanks soo much:)

OpenStudy (anonymous):

lol, glad to be of help :)